# Continual Imitation Learning for Prosthetic Limbs

### v1

@@ III-D Summary of proposed approach
$\mathcal{L}_{g}=\sum_{k=1}^{M_{t}-1}\left\lVert\mathbf{x}_{t,k+1}-g_{t}(\mathbf{x}_{t,k},{y}_{t,k})\right\rVert^{2}$ (1)
@@
$\hat{\mathbf{x}}_{t^{\prime},k}$ $=g_{t^{\prime}}\big{(}\mathbf{x}_{t^{\prime},k-1},$
$f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k-T-1},...,\mathbf{x}_{t^{\prime},k-1})]\big{)}$
@@
Our overall approach is shown in Algorithm 1 and Figures 1 and 2. We add new tasks $t$ one by one and split the training data for task $t$ into a training and a validation set. We train a newly evolved task-specific layer $f_{t}$ and the shared backbone $f^{s}$ to minimize the distance between $f_{t}[f^{s}(\mathbf{x}_{t,k-T}...\mathbf{x}_{t,k})]$ and ${y}_{t,k}$, and, at the same time, we train the shared backbone $f^{s}$ and the old task layers $f_{t^{\prime}}$ to minimize the distance between $f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k-T}...\mathbf{x}_{t^{\prime},k-1},\hat{\mathbf{x}}_{t^{\prime},k})]$ and ${y}_{t^{\prime},k}$ for the data in the rehearsal buffer. The rehearsal buffer contains both the original validation samples $\mathbf{x}_{t^{\prime},k}$, and the prospective samples $\hat{\mathbf{x}}_{t^{\prime},k}$ for a subsample $V_{t^{\prime}}$ of the time steps $k$ in the validation set. The prospective samples are generated via a task-specific, prospective model $g_{t^{\prime}}$ as $\hat{\mathbf{x}}_{t^{\prime},k}=g_{t^{\prime}}\big{(}\mathbf{x}_{t^{\prime},k-1},f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k-T-1}...\mathbf{x}_{t^{\prime},k-1})]\big{)}$.
IV Experiments
Datasets and metrics. We use three real-world human gait experiments’ datasets. The first two, ENABL3S and Embry [38, 39], are publicly available from popular gait labs, with ENABL3S providing approximately 5.2M training and 1.3M test samples from ten subjects across various locomotion activities, and Embry offering 512K training and 104K test samples from ten subjects also covering a range of movements. The third dataset is a novel collection from our lab, focusing on gait patterns of patients (transtibial amputees), recorded with a 200 Hz infrared motion capture system using twelve cameras, and includes activities like walking, stair ascent, and descent, with around 27K training and 11K testing samples. Ethical clearance for these experiments was obtained from our institutional review board. This dataset, along with the others, includes motion-related variables such as 3D angles of body joints and segments, velocities, and accelerations. To evaluate the accuracy of the desired joint profiles, $y_{k}$, we calculate the coefficient of determination ($R^{2}$), between the ground truth trajectories and model-based joint profiles.
Baselines. We conduct a comprehensive benchmarking of our approach against a range of learning-based limb behavior modeling methods documented in the literature, ranging from traditional linear models and Linear SVR, through advanced nonlinear models like SVR with RBF kernel, Random Forest (RF), and Decision Trees (DT), to neural networks including FFAN and RNNs with LSTM and GRU. Our model outperforms established benchmarks, providing consistent performance across various datasets (Figure 3).
Figure 3: Evaluation of our proposed multitask prospective rehearsal-integrated model across three distinct datasets (ENABL3, Embry et al., Amputees) in comparison with established baselines cited in the literature
Datasets $\rightarrow$ ENABL3S
Embry et. al. 2018
Speed ($ms^{-}1$) / Incline (${}^{\text{o}}$)
@@ IV Experiments
Tasks $\rightarrow$
Level
Walk
Ramp
Ascent
Ramp
Descent
Stair
Ascent
Stair
Descent
0.8/-10 0.8/0 0.8/10 1.0/-10 1.0/0 1.0/10 1.2/-10 1.2/0 1.2/10
Level
Walk
Stair
Ascent
Stair
Descent
head rehearsal Joint multi-task learning
@@
head rehearsal Task-incremental learning
@@
TABLE I: Ablation results: $R^{2}$ values for comparing rehearsal methods and architectures in joint and continual (task-incremental) settings.
Ablations. We assess the contribution of each component on overall performance. We explore both joint and continual learning paradigms, comparing models with shared final layers against those with task-specific final layers. Further, we consider both conventional and prospective rehearsal with each of the shared and task-specific final layer variants. In joint training, the training data is either augmented with the proposed multitask prospective rehearsal from a validation set or simply with the original samples from the validation set itself (conventional rehearsal). Note that, in all conditions, we used the same amount of training data to maintain fairness. Augmentation with prospective rehearsal improved the performance of models with both shared and task-specific final layers. Further, the results provide evidence that task-specific final layers help to account for particularities among tasks. A comparison of performance in task-incremental learning yields similar results. The results demonstrate that the choice of a combination of prospective rehearsal and an evolving architecture, where task-specific final layers are integrated with a shared backbone, yields the best performance across the board (Table I). All following experiments are performed with task-specific final layers.
Dataset ENABL3S
Embry et. al. 2018
Speed ($ms^{-}1$)/Incline(${}^{\text{o}}$)
Transtibial
amputees
Task Level Walk Ramp Ascent Ramp Descent Stair Ascent Stair Descent 0.8/-10 0.8/0 0.8/10 1.0/-10 1.0/0 1.0/10 1.2/-10 1.2/0 1.2/10 Level Walk Stair Ascent Stair Descent
@@
TABLE II: Comparison of performance ($R^{2}$) of proposed prospective rehearsal with other representative continual learning strategies – SI (Synaptic Intelligence), EWC (Elastic weight consolidation), PNN (Progressive Neural Networks), GEM (Gradient episodic memory), ER (Experience Replay).
Datasets ENABL3S
Embry et. al. 2018
Speed ($ms^{-}1$) / Incline (${}^{\text{o}}$)
@@
Tasks
Level
Walk
Ramp
Ascent
Ramp
Descent
Stair
Ascent
Stair
Descent
0.8/-10 0.8/0 0.8/10 1.0/-10 1.0/0 1.0/10 1.2/-10 1.2/0 1.2/-10
Level
Walk
Stair
Ascent
Stair
Descent
TABLE III: Comparison of performance ($R^{2}$) of Gaussian noise injection and proposed prospective rehearsal.
Benchmarking continual learning strategies.
In the continual learning paradigm, we compare our approach against regularization strategies like SI and EWC, architectural strategies, PNNs, and state-of-the-art replay-based strategies like GEM and the classic form of experience replay, which we refer to as conventional rehearsal (Table II). We observe that prospective rehearsal performs best across the board. In Table III, we compare data augmentation via Gaussian noise injection to the proposed prospective rehearsal, with prospective rehearsal performing best across the board.
Resilience to distribution shift. In some cases, we do not observe a significant improvement in the performance of the prospective rehearsal over a conventional rehearsal. To explain these results, we inspect the difference in $R^{2}$ between the prospective and conventional rehearsal versus the Jenson-Shannon (JS) distance [40] between training and test distributions (Figure 4). We observe that the error difference positively correlates with JS distance. In other words, the advantage of the prospective rehearsal becomes more pronounced the more training and test distribution deviate. This is in line with our theoretical analysis: The more training and test distribution differ, the more prediction errors of $f$ we expect, which means that our proposed prospective rehearsal helps more. Since real-world prosthesis data is prone to noise and variability from various exteroceptive and interoceptive factors [41], being robust to distribution shifts is an important requirement for this domain, and our approach performs well in this regard.
Figure 5: A) Performance comparison of prospective and conventional rehearsal strategies across different strengths, $\tau$, of adversarial perturbations. B) Performance of linear and non-linear probing on downstream locomotion task classification. Prospective rehearsal training performs best across the board.
@@
Figure 6: (A) (Left) Results of statistical significance tests comparing the performance of conventional and proposed prospective rehearsal for shared and task-specific final layers. (Right) Comparison of performance of the proposed prospective rehearsal with state-of-the-art continual learning strategies. (B) Comparison of the performance of continually trained models with conventional and proposed prospective rehearsal for different levels of input noise. Results are shown for models with task-specific final layers. (C) Plot depicting NRMSE values for test tasks from Embry and ENABL3S datasets post-training, with the X-axis for learned tasks, color for test tasks, and shading indicating standard error across subjects. Statistical significance was assessed using the Wilcoxon-signed rank test with Bonferroni correction, ${}^{****}:p<1e-4,^{***}:p<1e-3,^{**}:p<1e-2$.
Resilience to adversarial perturbations. To further assess the resilience of the prospective rehearsal approach, we compare its outcomes to those without it in the face of adversarial perturbations. These adversarial samples are crafted using the Fast Gradient Sign Method (FGSM) [42] across diverse epsilon thresholds, $\tau$. Notably, our proposed prospective rehearsal technique surpasses the conventional baseline across all evaluated FGSM attack intensities (Figure 5A). Furthermore, as the perturbation intensity escalates, the performance disparity between the prospective and the conventional method widens, reinforcing the superior adaptability of prospective rehearsal under scenarios where test data distribution deviates from the training set. The obtained results are statistically significant.
Probing downstream performance. We evaluate the representation capacity of models trained with prospective rehearsal techniques versus those trained without it, focusing on a downstream challenge—classifying different types of locomotion. We employ both linear and nonlinear (MLP) probes to assess the representation of the shared layers, denoted as $f_{s}(\textbf{x}_{t,k-T}...\textbf{x}_{t,k})$, from models with task-specific heads aimed at classifying the task $t$. The findings indicate that prospective rehearsal yields significantly better task classification performance than its conventional counterpart for both linear and nonlinear probes (Figure 5B). This result underscores that the prospective model-based training confers an added benefit of robustness and adaptability compared to conventional rehearsal techniques.
Resilience to noise. Further, we analyze the robustness of our proposed model against perturbations to the input signals during test time in the form of multivariate white Gaussian noise, $\mathcal{N}(\mathbf{0},\mathbf{\Sigma}\in\mathbb{R}^{d\times d})$. We apply different levels of perturbation $L$, where $L$ is defined as the ratio of the standard deviation of the Gaussian noise to the standard deviation of the test inputs. Figure 6B shows the $R^{2}$ of our proposed rehearsal and a conventional rehearsal for five different tasks on the ENABL3S dataset. While performance generally decreases with higher amounts of noise, the prospective rehearsal yields better $R^{2}$ across the board.
@@
Stability against forgetting. We explore the ability to preserve the knowledge of previously learned tasks after learning new ones. Figure 6C shows how the NRMSE values (lower is better) for each task develop when training subsequent tasks, both for a shared final layer (top) and for a task-specific final layer (bottom). Both architectures preserve the memory of the previous tasks, but the performance on old tasks slightly worsens for a shared final layer, whereas it remains the same for task-specific final layers.

### v3

@@ 3 Method
$\mathcal{L}_{g}=\sum_{k=1}^{M_{t}-1}\left\lVert\mathbf{x}_{t,k+1}-g_{t}(\mathbf{x}_{t,k},{y}_{t,k})\right\rVert^{2}$
@@
                   $\hat{\mathbf{x}}_{t^{\prime},k}=g_{t^{\prime}}\big{(}\mathbf{x}_{t^{\prime},k-1},f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k-T-1},...,\mathbf{x}_{t^{\prime},k-1})]\big{)}$
@@
4 Experiments
Datasets and metrics. We use three real-world human gait experiments’ datasets. The first two, ENABL3S (Hu et al., 2018) and Embry (Embry et al., 2018), are publicly available from popular gait labs. ENABL3S comprises approximately 5.2 million training samples and 1.3 million test samples collected from ten subjects performing various locomotion activities, including level-ground walking, stair ascent and descent, and ambulating on an inclined walkway. From the Embry dataset, we selected data representing nine very distinct combinations of walking speeds (ranging from 0.8 m/s to 1.2 m/s) and inclines (ranging from -10 degrees to +10 degrees). This selection resulted in approximately 512,000 training samples and 104,000 test samples from ten subjects, encompassing a diverse range of movements. The third dataset is a novel collection from our lab, focusing on gait patterns of patients (transtibial amputees), recorded with a 200 Hz infrared motion capture system using twelve cameras, and includes activities like walking, stair ascent, and descent, with around 27K training and 11K testing samples. Ethical clearance for these experiments was obtained from our institutional review board. This dataset, along with the others, includes motion-related variables such as 3D angles of body joints and segments, velocities, and accelerations. To evaluate the accuracy of the desired joint profiles, $y_{k}$, we calculate the coefficient of determination ($R^{2}$), between the ground truth trajectories and model-based joint profiles.
Table 1: Experimental results for (A) joint multitask learning using different prediction head architectures and replays, (B) task incremental learning using different prediction head architectures and replays, (C) different continual learning strategies – SI (Synaptic Intelligence), EWC (Elastic weight consolidation), PNN (Progressive Neural Networks), GEM (Gradient episodic memory), ER (Experience Replay), and proposed prospective replay, (D) data augmentation with Gaussian noise vs. prospective replay buffer. (LW=Level walking, RA=Ramp ascent, RD=Ramp descent, SA=Stair ascent, SD=Stair descent, Spd=Speed in m/s, Inc=Inclination in °)
Dataset ENABL3S
Embry
et. al. 2018
Transtibial
@@ 4 Experiments
Task LW RA RD SA SD Spd 0.8 0.8 0.8 1.0 1.0 1.0 1.2 1.2 1.2 LW SA SD
Inc -10 0 10 -10 0 10 -10 0 10
head replay (A) Joint multi-task learning
@@
head replay (B) Task-incremental learning
@@
(C) Continual learning strategies
@@
(D) Data augmentation
Ablations. We assess the contribution of each component on overall performance. We explore both joint and continual learning paradigms, comparing models with shared final layers against those with task-specific final layers Dey et al. (2024). Further, we consider both conventional and prospective rehearsal with each of the shared and task-specific final layer variants. In joint training, the training data is either augmented with the multitask prospective rehearsal from a validation set or simply with the original samples from the validation set itself (conventional rehearsal). In all conditions, we used the same amount of training data to maintain fairness. Augmentation with prospective rehearsal improved the performance of models with both shared and task-specific final layers. Further, the results provide evidence that task-specific final layers help to account for particularities among tasks. A comparison of performance in task-incremental learning yields similar results. The results demonstrate that the choice of a combination of prospective rehearsal and an evolving architecture, where task-specific final layers are integrated with a shared backbone, yields the best performance across the board (Table 1A–B). All following experiments are performed with task-specific final layers.
Benchmarking continual learning strategies. In the continual learning paradigm, we compare our approach against regularization strategies like SI and EWC, architectural strategies, PNNs, and state-of-the-art replay-based strategies like GEM and the classic form of experience replay, which we refer to as conventional rehearsal (Table 1C). We observe that prospective rehearsal performs best across the board, indicating better task memory retention and generalization. In Table 1D, we compare data augmentation via Gaussian noise injection to the proposed prospective rehearsal, with prospective rehearsal performing best across the board. All results are statistically significant (see appendix).
Resilience to distribution shift. In some cases, we do not observe a significant improvement in the performance of the prospective rehearsal over a conventional rehearsal. To explain these results, we inspect the difference in $R^{2}$ between the prospective and conventional rehearsal versus the Jenson-Shannon (JS) distance (Endres & Schindelin, 2003) between training and test distributions (Figure 3). We observe that the error difference positively correlates with JS distance. In other words, the advantage of the prospective rehearsal becomes more pronounced the more training and test distribution deviate. This is in line with our theoretical analysis: The more training and test distribution differ, the more prediction errors of $f$ we expect, which means that our proposed prospective rehearsal helps more. Since real-world prosthesis data is prone to noise and variability from various exteroceptive and interoceptive factors (Prahm et al., 2019), being robust to distribution shifts is an important requirement for this domain, and our approach performs well in this regard.
Figure 4: A) Performance comparison of prospective and conventional rehearsal strategies across different strengths, $\tau$, of adversarial perturbations. B) Performance of linear and non-linear probing on downstream locomotion task classification. Prospective rehearsal training performs best across the board. (C) Comparison of the performance of continual models with conventional and prospective rehearsal for different levels of input noise. Results are presented for models with task-specific final layers.
Resilience to adversarial perturbations. To further assess the resilience of the prospective rehearsal approach, we compare its outcomes to those without it in the face of adversarial perturbations. These adversarial samples are created using the Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014) across diverse epsilon thresholds, $\tau$. Notably, our prospective rehearsal technique surpasses the conventional baseline across all evaluated FGSM attack intensities (Figure 4A). Furthermore, as the perturbation intensity escalates, the performance disparity between the prospective and the conventional method widens, reinforcing the superior adaptability of prospective rehearsal under scenarios where test data distribution deviates from the training set. The results are statistically significant.
Probing downstream performance. We evaluate the representation capacity of models trained with prospective rehearsal techniques versus those trained without it, focusing on a downstream challenge—classifying different types of locomotion. We employ both linear and nonlinear (MLP) probes to assess the representation of the shared layers, denoted as $f_{s}(\textbf{x}_{t,k-T}...\textbf{x}_{t,k})$, from models with task-specific heads aimed at classifying the task $t$. The findings indicate that prospective rehearsal yields significantly better task classification performance than its conventional counterpart for both linear and nonlinear probes (Figure 4B). This result underscores that the prospective model-based training confers an added benefit of robustness and adaptability compared to conventional rehearsal.
Resilience to noise. Further, we analyze the robustness of our proposed model against perturbations to the input signals during test time in the form of multivariate white Gaussian noise, $\mathcal{N}(\mathbf{0},\mathbf{\Sigma}\in\mathbb{R}^{d\times d})$. We apply different levels of perturbation $L$, where $L$ is defined as the ratio of the standard deviation of the Gaussian noise to the standard deviation of the test inputs. Figure 4C shows the $R^{2}$ of our proposed rehearsal and a conventional rehearsal for five different tasks on the ENABL3S dataset. While performance generally decreases with higher amounts of noise, the prospective rehearsal yields better $R^{2}$ across the board.
Figure 5: (Left) Snapshots from the experiments on a subject. Different motion conditions such as walking at different speeds and inclines were tested on an instrumented treadmill with adjustable inclines (A–E). Furthermore, ground walking on a level walkway with different motion maneuvers was also tested (F–N). (Right) Stability comparison of foot kinematic profiles yielded using prospective and conventional rehearsals. First $\lambda_{1}$ and second $\lambda_{2}$ Lyapunov exponents (lower is better), computed using the Eckmann method, indicate that the foot kinematic profiles obtained using our proposed method exhibit greater stability compared to those obtained without employing our method. Median Lyapunov exponent of the prospective rehearsal method, (dotted line) is lower compared to conventional rehearsal (solid line).
Gait stability assessment with an exoskeleton. In a pilot study, able-bodied individuals walked with an ankle exoskeleton (Figure 5 left) controlled by joint profiles predicted by the model with the prospective rehearsal approach. Participants reported that the predictions felt timely, accurate, and in sync with their residual body movements. To evaluate the stability of the model-based foot profiles, with and without the prospective rehearsal approach, we measured the maximal Lyapunov exponent using the Eckmann method (Eckmann et al., 1986). The Lyapunov exponent measures the rate of divergence of nearby trajectories in a dynamical system, with lower values indicating higher stability. Our results showed that both the first and second Lyapunov exponents were lower when using our proposed method, indicating more stable gait trajectories (Figure 5 right).
5 Conclusion
Author Contributions
Sharmita Dey led the project, conceptualized and developed the idea, designed the algorithm, developed the frameworks, conducted computational and hardware experiments, and wrote and revised the manuscript. Benjamin Paassen contributed to the theoretical framework and revised the manuscript. Sarath Ravindran Nair contributed to computational experiments and revised the manuscript. Sabri Boghourbel contributed to important discussions and revision of the manuscript. Arndt Schilling contributed to important discussions and revision of the manuscript.
Appendix A Additional Results
A.1 Model architecture
In Table 2, we show the results of using different architectures for the shared backbone $f^{s}$. All models were designed to have approximately the same number of parameters. Since TCN tends to perform best across datasets, we opt for TCN as the backbone architecture in practice.
Dataset ENABL3S
Embry
et. al. 2018
@@ A.1 Model architecture
Task LW RA RD SA SD Spd 0.8 0.8 0.8 1.0 1.0 1.0 1.2 1.2 1.2 LW SA SD
Inc -10 0 10 -10 0 10 -10 0 10
Linear 0.68 0.89 0.89 0.90 0.83 0.04 0.15 0.51 0.04 0.12 0.55 0.13 0.11 0.55 0.39 0.64 0.68
MLP 0.67 0.91 0.91 0.91 0.86 -0.06 0.14 0.52 0.06 0.13 0.55 0.11 0.12 0.56 0.42 0.56 0.70
LSTM 0.88 0.94 0.94 0.96 0.93 0.86 0.97 0.95 0.83 0.96 0.96 0.82 0.95 0.97 0.41 0.59 0.64
Transformer 0.9 0.94 0.95 0.96 0.93 0.88 0.97 0.96 0.86 0.98 0.97 0.88 0.97 0.98 0.98 0.84 0.92
TCN 0.91 0.95 0.95 0.97 0.94 0.96 0.98 0.95 0.97 0.99 0.97 0.91 0.97 0.98 0.98 0.91 0.92
Table 2: Comparison of various shared module architectures on different locomotion conditions from three datasets.
A.2 Model predictions
In Figure 6 (left), We show the model-based motion trajectories of a transtibial amputee subject’s hypothetical ankle, as demonstrated by our multitask model using prospective rehearsal. The alignment between the predicted and biological able-bodied limb motions suggests the model’s effectiveness in approximating gait synergy across various lower limb joints and segments, aiming to replicate normal locomotion for impaired patients. Figure 6 (right) presents analogous results for an able-bodied subject, using their own limb motion as the target. The close match between model-based and natural limb motions further underscores the model’s potential in generating movement signals for bionic limbs, offering hope for enhancing mobility in individuals with lower-extremity disabilities.
Figure 6: Model-based kinematic profiles from our proposed rehearsal-based models. (Left) Motion trajectories of a transtibial amputee subject’s hypothetical ankle, predicted by our model. The normative (target) trajectories are computed from an able-bodied individual with anthropometric features and walking speed similar to that of the amputee subject. (Right) Predicted motion trajectories of an able-bodied subject from ENABL3S dataset for different locomotion modes. The target trajectories in this case are the subject’s own limb motion. In both cases, the predicted trajectories are highly aligned with the target limb motion.
In Figure 7 we show the example predictions from baseline models as compared to our approach. The joint angular positions predicted by the baseline models (shared final layer model and EWC regularization) showed frequent misalignments with sharp disagreements between the predicted and actual trajectories. In a safety-critical scenario such as prosthesis control for locomotion assistance, such misalignments can be very costly as they can lead to imbalance. On the other hand, predictions from our proposed model show a higher degree of alignment with the ground truth (actual) trajectories, which is necessary to ensure smoother locomotion using a prosthetic device controlled by these predictions.
Figure 7: Baseline prediction examples. (Left) Examples of joint position trajectories for different locomotion tasks predicted by a shared final layer model vs. task-specific final layer model. (Right) Examples of joint position trajectories predicted by a task-specific final layer model with EWC regularization vs. task-specific final layer model with prospective rehearsal. The black curve shows the ground truth trajectories.
A.3 Resilience to distribution shift
Figure 8: (Left) Train and test distributions for two tasks from Embry et. al. dataset for a representative subject. (Right) Predictions of model trained with conventional and prospective rehearsal.
We show the effectiveness of our prospective rehearsal-based approach in dealing with distribution shifts during test time from the lens of model predictions. Figure 8 illustrates the predictions of models trained with conventional rehearsal and our prospective rehearsal for two scenarios—one with a low and the other with a high distribution shift during test time. As anticipated, both models exhibit comparable performance when the test and training distributions are similar (see the bottom of Figure 8). However, in the presence of a distribution shift at test time, the model trained with conventional rehearsal shows a decline in performance, while the prospective rehearsal-trained model sustains its predictive accuracy (see the top of Figure 8).
A.4 Stability against forgetting
@@
Figure 9: (A) Plot depicting NRMSE values for test tasks from Embry and ENABL3S datasets post-training, with the X-axis for learned tasks, color for test tasks, and shading indicating standard error across subjects. (B) (Left) Results of statistical significance tests comparing the performance of conventional and proposed prospective rehearsal for shared and task-specific final layers. (Right) Comparison of performance of the proposed prospective rehearsal with state-of-the-art continual learning strategies. Statistical significance was assessed using the Wilcoxon-signed rank test with Bonferroni correction, ${}^{****}:p<1e-4,^{***}:p<1e-3,^{**}:p<1e-2$.
@@
We explore the ability to preserve the knowledge of previously learned tasks after learning new ones. Figure 9A shows how the NRMSE values (lower is better) for each task develop when training subsequent tasks, both for a shared final layer (top) and for a task-specific final layer (bottom). Both architectures preserve the memory of the previous tasks, but the performance on old tasks slightly worsens for a shared final layer, whereas it remains the same for task-specific final layers.
Train task
Walk Ramp up Ramp down Stair up Stair down
Test task Walk 0.054 0.14 0.16 0.177 0.18
Ramp up 0.454 0.055 0.147 0.134 0.176
Ramp down 0.517 0.517 0.055 0.165 0.156
Stair up 0.516 0.516 0.516 0.054 0.193
Stair down 0.506 0.506 0.502 0.505 0.066
Table 3: RMSE values of a model with task-specific prediction heads trained without replay on different tasks after it is trained for a particular task.
Train task
Walk Ramp up Ramp down Stair up Stair down
Test task Walk 0.054 0.062 0.067 0.071 0.078
Ramp up 0.15 0.06 0.067 0.074 0.075
Ramp down 0.175 0.19 0.064 0.083 0.083
Stair up 0.213 0.184 0.215 0.071 0.083
Stair down 0.235 0.231 0.204 0.228 0.086
Table 4: RMSE values of a model with task-shared prediction heads trained with replay on different tasks after it is trained for a particular task.
Train task
Walk Ramp up Ramp down Stair up Stair down
Test task Walk 0.053 0.057 0.058 0.059 0.062
Ramp up 0.454 0.057 0.055 0.056 0.056
Ramp down 0.502 0.501 0.051 0.056 0.057
Stair up 0.553 0.551 0.551 0.053 0.054
Stair down 0.53 0.53 0.53 0.53 0.066
Table 5: RMSE values of a model with task-specific prediction heads trained with replay on different tasks after it is trained for a particular task.
We further show through ablation studies that, both the rehearsal and architectural choice are necessary for preventing the forgetting problem (see Tables 35). A model trained without replay (Tab. 3) excels only on its current task, but suffers from catastrophic forgetting on previous tasks. In contrast, models with replay (Tab. 4 and 5) reduce forgetting. Furthermore, using task-specific prediction heads outperforms a shared head, further mitigating forgetting. Thus, both replay and task-specific heads help address forgetting.
In addition, to systematically analyze the model performance as it is trained for multiple tasks, we also computed two metrics (similar to Díaz-Rodríguez et al. (2018)) to measure how well the model performs on previously trained tasks after training it with new tasks.
The backward transfer is computed as
$\mathrm{BWT}(t)=\frac{1}{t-1}\sum_{j=1}^{t-1}\varepsilon_{j}(j)-\varepsilon_{t}(j),$ (1)
where $t$ is the latest task the model is trained for, $\varepsilon_{i}(j)$ is the prediction error on task $j$ once the model was trained with the task $i$. $\mathrm{BWT}(t)$ computes the change in performance of the model on previously trained tasks, $j=1..t-1$ after it is trained for a new task $t=2..T$. Larger values of backward transfer are preferred.
The forgetting ratio for a task $t$ is computed as
$\mathrm{FR}(t)=\frac{\varepsilon_{T}(t)-\varepsilon_{t}(t)}{\varepsilon_{t}(t)},$ (2)
where $T$ represents the final task on which the model was trained. $\mathrm{FR}(t)$ gives the relative change in performance of the model on task, $t=1..T-1$, after it was trained with all available tasks. Smaller values of the forgetting ratio are preferred.
Train task, $t$ Walk Ramp up Ramp down Stair up Stair down
no replay -0.043 -0.065 -0.079 -0.098
task-shared with replay -0.004 -0.006 -0.013 -0.014
task-specific with replay -0.002 -0.001 -0.002 -0.003
Table 6: Backward knowledge transfer, $\mathrm{BWT}(t)$, (transfer of knowledge to previously learned tasks) when the model is trained with a new task, $t$. Higher the better.
Evaluation task, $t$ Walk Ramp up Ramp down Stair up Stair down
no replay 2.33 2.29 1.99 2.64
task-shared with replay 0.42 0.25 0.27 0.16
task-specific with replay 0.16 0 0.1 -0.02
Table 7: Forgetting ratios, $\mathrm{FR}(t)$, of the model on a task, t after training it with all available tasks. The lower the better.
Tables 6 and 7 show the backward transfer and forgetting ratios of models trained without replay, and the ones trained with replay with different architectures. These results further illustrate that the models with task-specific prediction heads and replay perform best in the face of continual task learning and dealing with catastrophic forgetting.
A.5 Statistical significance of results